1-Calculate the third-degree Taylor polynomial about x0=0 for f (x) =(1+ x)1/2

2-Use the polynomial in part (i) to approximate (1.1)^1/2 and find a bound for the

error involved.

3.Use the polynomial in part (i) to approximate integration[0,0.1]

(1 +x)1/2dx .



1
Expert's answer
2021-05-07T14:32:00-0400

1.

"f(x)=(1+x)^{\frac{1}{2}} \newline
f'(x)=\frac{1}{2}(1+x)^{-\frac{1}{2}}\newline
f''(x)=-\frac{1}{4}(1+x)^{-\frac{3}{2}}\newline
f'''(x)=\frac{3}{8}(1+x)^{-\frac{5}{2}}\newline
Taylor series,\newline
f(x)=\sum_{n=0}^{3}\frac{f^n(x_{0})}{n!}(x-x_{0})^n\newline
f(x)=f(x_{0})+\frac{f'(x_{0})}{1!}(x-x_{0})+\frac{f''(x_{0})}{2!}(x-x_{0})^2+\frac{f'''(x_{0})}{3!}(x-x_{0})^3\newline
f(x)=f(0)+\frac{f'(0)}{1!}(x-0)+\frac{f''(0)}{2!}(x-0)^2+\frac{f'''(0)}{3!}(x-0)^3\newline
f(x)=f(0)+\frac{f'(0)}{1!}(0)+\frac{f''(0)}{2!}(0)^2+\frac{f'''(0)}{3!}(0)^3\newline
f(x)=1+\frac{1}{2}x-\frac{1}{8}x^2+\frac{3}{48}x^3"


2.

Given, (1.1)1/2.

Actual value of (1.1)1/2=1.048808848. (9 decimal places)

"f(0.1)=(1.1)^{\frac{1}{2}} \newline
f(0.1)=1+\frac{1}{2}(0.1)-\frac{1}{8}(0.1)^2+\frac{3}{48}(0.1)^3\newline
f(0.1)=1.048812500"

Error="\begin{vmatrix}
 approx -exact \\

\end{vmatrix}
=\begin{vmatrix}
 1.0488125 -1.048808848 \\

\end{vmatrix}
=0.000003652"


3.

Approximate "\int_{0}^{0.1}(1+x)^\frac{1}{2}dx" ,


"\int_{0}^{0.1}f(x)dx=\int_{0}^{0.1}(1+\frac{1}{2}x-\frac{1}{8}x^2+\frac{3}{48}x^3)dx=0.102459896"